# Training–free AI for Earth Observation Change Detection using Physics Aware Neuromorphic Networks

### v1

@@ Abstract
Earth observations from low Earth orbit satellites provide vital information for decision makers to better manage time–sensitive events such as natural disasters. For the data to be most effective for first responders, low latency is required between data capture and its arrival to decision makers. A major bottleneck is in the bandwidth–limited downlinking of the data from satellites to ground stations. One approach to overcome this challenge is to process at least some of the data on-board and prioritise pertinent data to be downlinked. In this work we propose a Physics Aware Neuromorphic Network (PANN) to detect changes caused by natural disasters from a sequence of multi-spectral satellite images and produce a change map, enabling relevant data to be prioritised for downlinking. The PANN used in this study is motivated by physical neural networks comprised of nano-electronic circuit elements known as “memristors” (nonlinear resistors with memory). The weights in the network are dynamic and update in response to varying input signals according to memristor equations of state and electrical circuit conservation laws. The PANN thus generates physics–constrained dynamical output features which are used to detect changes in a natural disaster detection task by applying a distance-based metric. Importantly, this makes the whole model training–free, allowing it to be implemented with minimal computing resources. The PANN was benchmarked against a state-of-the-art AI model and achieved comparable or better results in each natural disaster category. It thus presents a promising solution to the challenge of resource–constrained on-board processing.
@@ Introduction
Deep ML models, specifically CNNs, have been applied to several natural disaster tasks using EO data, including flood segmentation [13, 14], detecting earthquake affected buildings [15] and volcanic eruptions [16]. To increase the performance of these models, often the number of layers and weights within the model needs to be increased. Chintalapati et al. [17] showed how the architecture of CNN models has evolved to increase performance on an EO image classification task, which resulted in an increase in the number of weights in the models, with the two best performing models, CoCa [18] and BASIC-L [19], using 2.1 and 2.44 billion parameters, respectively. It was noted that most deep learning models have so many parameters that they would exceed the memory budget on most EO satellites. In recent years, work has been done to design smaller AI models specifically for use on-board satellites [20, 21]. Furthermore, missions such as $\Phi$-Sat-1[22] and OPS-SAT [23] have also designed satellites to accommodate AI models on-board. Although these models have been specifically designed for on-board use, all the training still occurs on the ground because of the large computational overhead. Only inference is performed on-board. Since AI models are trained on the ground, they often cannot be trained with the on-board sensor data, leading to degraded performance [24, 25]. Work has been done to retrain models, once operational, with new data as it is acquired [26, 27].
@@
In this study, we introduce the PANN, a AI model with a novel architecture modelled after physical neural networks comprised of nano-electronic circuit elements known as memristors (nonlinear resistors with memory) that mimic synapses [28, 29]. Like biological synapses, the memristive network weights continuously update in response to varying input signals, constrained by physical laws embedded within the model. The PANN, therefore, actively learns as new inputs are fed into the network, such as from a sensor on-board a satellite. As such, the PANN takes a significant shift away from the conventional artificial neural network AI paradigm, as the gradient–descent, back–propagation training of weights is completely replaced by physics–constrained dynamical equations. This is also in contrast to physics–informed neural networks, where the physics is embedded as a component of the model in the artificial neural network paradigm. Typically, this is done by modifying activation functions, gradient descent optimisation techniques, network structure, or loss functions [30, 31]. The added physics normally takes the form of a partial differential equation and is task specific, for example, incompressible Navier-Stokes equations for fluid flow problems [32], whereas with the PANN, the physics is independent of the task.
@@
In this study, we use a natural disaster dataset released by Růžička et al. [21] to evaluate the performance of the PANN model on a change detection problem, that contains a time series of multi-spectral EO images from Sentinel-2. We benchmark the PANN against a state-of-the-art deep learning model and achieve comparable or better results. Additionally, we use a distance measurement to detect changes, bypassing the need for an output layer altogether. Hence, we do not need to train any part of the model and the PANN does not need to be trained on the ground prior to deployment.
@@ Results
The PANN model is used here to detect regions of change from natural disasters using EO multi-spectral images. The change detection task requires the model to take in a sequence of five images, where the final image is after a natural disaster event. The model outputs a change feature map that highlights the regions affected by the natural disaster, by breaking the images into smaller tiles and assigning a change likelihood score for each tile. The PANN model used in this study is an adaptive network, with weights that adjust with each new input. This allows the PANN to generate features from the inputs without requiring any training (see Methods for details). To determine if an image area has changed, the distance between inputs are compared in the feature space rather than the input space (pixel comparison). By applying a distance-based metric directly to the feature space, no training is performed at any stage in the model. The dataset includes four different categories of natural disaster events: fires, floods, hurricanes and landslides. Additionally, each natural disaster event is accompanied by a change mask which is used only to evaluate the performance of the PANN model.
@@
Figure 1 shows the output change maps produced by the PANN model for each type of natural disaster event. Alongside the change maps are the RGB images directly before and after the natural disaster events and the associated change mask, which also includes clouds from the images directly before and after the natural disaster event. Although only one image is shown before the natural disaster event in Fig. 1, a total of four images before the natural disaster event were used to create the displayed change map. For the four natural disaster events, the main areas of change have been correctly identified as changes in the change maps, with a limited amount of noise in the images being considered as a change, such as the vegetation colour change in the fire natural disaster event (top row), or the sediment being washed into the ocean in the flood natural disaster event (second row). Clouds are detected as changes, but are masked out of the change maps shown in Fig. 1, since they are ignored when evaluating the change maps to allow for a more direct comparison to a state-of-the-art deep learning model (see Methods for details).
@@
Figure 1: RGB images (columns 1 and 2) directly before and after a natural disaster event, the corresponding target change mask (column 3) and the output change map produced by the PANN model (column 4) for each type of natural disaster: Top row – an example fire natural disaster event; Second row – an example flood natural disaster event; Third row – an example hurricane natural disaster event; Bottom row – an example landslide natural disaster event. Cloud-cover has been masked from the images and corresponds to yellow in the change mask, with genuine differences coloured aqua.
@@
For each natural disaster event a change map was created, with the order in which the natural disasters were passed into the PANN model being randomised. We evaluated the quality of the change maps by calculating the Area Under the Precision—Recall Curve (AUPRC). The change likelihood score for each tile is given by equation (3) in Methods, which depends on the distance metric used to measure the position of the tiles in the feature space. Table 1 reports the overall AUPRC as a percentage for each of the natural disaster categories using three different distance metrics, Euclidean, Cosine and Correlation distances. The Correlation distance significantly outperforms the other distance metrics and is also much more consistent across the different natural disaster classes.
@@
Table 1: Comparison between different distance metrics used to calculate the PANN change score for different natural disaster classes. All values are calculated as mean AUPRC (%) $\pm$±SEM over 5 runs. Best scores are highlighted.
@@
Table 2: Comparison between change scores predicted by different models for different natural disaster classes. All values are calculated as AUPRC (%), with values for the PANN model corresponding to the best scores from Table 1.
@@
The AI models RaVAEn and PANN outperform the baseline method, since they can extract the most relevant features from the tiles and group similar tiles together, while keeping dissimilar tiles further apart in their feature spaces. To visualise the high dimensional feature space of the PANN, we used the Uniform Manifold Approximation and Projection (UMAP) [50] method to reduce the dimensions of the feature space into 2D. The UMAP used 20 neighbours, a minimum distance of 0.4 and the correlation distance, unless otherwise stated. Figure 2 shows the UMAP 2D projection of the feature space for a flooding event. The left-hand side shows the position of the tiles in the image directly before and after the flooding event. For tiles after the flooding event, the percentage change in each tile is denoted by the colour bar. The right-hand side of the figure shows the same projection of the feature space but with the tile images instead of points. The change tiles are grouped together at the bottom of the projections and the remaining tiles are also sorted according to the features in the tiles, with greener farmland tiles on the left-hand side and tiles with the white objects on the right. Tiles that have missing values, from the north alignment of the swath, are also grouped towards the bottom right, with the tiles that have the missing values across the top being separated from the tiles with the missing values on the right-side of the tile. For comparison with the UMAP projection for the pixel space, see Supplementary Fig. 8.
@@ Discussion
In this study, we introduced the PANN model for a natural disaster change detection task and compared it against a state-of-the-art AI model RaVAEn and a baseline method. All three models receive the same multi-spectral EO images for the inputs and use the same change detection equation given by equation (3), for each tile in the image. The difference in the model performances, therefore, is attributed to their ability to extract the most relevant features from the inputs, while ignoring noise within the inputs. We note, however, that while the PANN model receives the same inputs as the other models, only the inputs corresponding to the bands selected by the automatic feature engineering process, on a scene by scene basis, are passed to the relevant individual networks. This inherently allows the PANN model to ignore irrelevant information as determined by the domain knowledge embedded in the feature engineering process. The baseline method does not extract any features from the inputs and instead does the comparison directly in the pixel space. This makes the baseline method more sensitive to noise in the images and, as expected, it exhibits the lowest performance as reported in Table 2. The machine learning models, on the other hand, extract meaningful features from the tiles and hence outperform the baseline method in every natural disaster category. The PANN and RaVAEn models achieve a comparable result in three of four of the natural disaster categories, demonstrating that both models extract equally useful features for these categories. For flood natural disasters, however, the PANN model outperforms the RaVAEn model and thus arguably extracts more meaningful features.
@@
While being able to extract the features is the most important task for the PANN model, how the feature space is navigated is also important for the performance on the PANN model. This is seen in Table 1 where the same feature space can produce significantly different results based on the distance metric used to navigate the feature space. This is visualised in the UMAP projections in Fig. 3, where different distance metrics are used to create the projections. The feature space produced by the PANN has a high dimensionality, as such the relative Euclidean distance between points tends to decrease[51]. This would result in the change likelihood score decreasing, particularly for change tiles and is likely why the Euclidean distance does not produce the best results. What is surprising is the large difference between the Cosine and Correlation distances, given that both metrics are angle-based measures. The Cosine distance is given by equation (4) and measures the angle between the vectors in the feature space. The Correlation distance is the same as the Cosine distance except each sample is mean centred first, as given by equation (5). The mean-correction for the Correlation distance effectively changes the location of the origin and hence the angle between the vectors also changes [52]. One possible reason why the mean correction has such a large impact on the performance is that it helps reduce covariate shift in the output sequence that can arise from the dynamical nature of the PANN. The covariate shift effect can be seen in Fig. 3 (and Supplementary Fig. 9), where there is a much greater overlap of points at the different timesteps (that are not changes) in the Correlation projection compared to that of the Cosine projection. Interestingly, batch normalisation layers were introduced to deep neural networks to help reduce the covariate shift between layers in the networks [53] and have been widely accepted to speed up training times and increase performance. We note that the PANN does not have any batch normalisation layers, while the RaVAEn model does, and that the best performing metrics for the models are the Correlation and Cosine distance, respectively.
@@
Our results show that the PANN model achieves comparable-to-better results than other state-of-the-art AI models for the natural disaster change detection task conducted in this study. The more impressive result here is that the PANN was able to achieve this result without any training to create the mapping from the input space to the feature space. The PANN is training-free as it continuously adapts with each new input. The model equations are relatively light–weight and thus, evaluation of the PANN model was able to be carried out using only CPUs in this study. The physics–based equations that constitute the PANN model are a significant deviation from conventional deep multi-layered bipartite neural networks that typically use mathematical activation functions to perform nonlinear transformations and stochastic gradient decent with backpropagation as the learning algorithm to train networks weights into an optimal static state. As such, the iterative training process is computationally expensive and usually requires cloud-based GPUs [54], to produce a static model.
@@
In conclusion, we demonstrated a training-free AI model, the PANN, for a natural disaster EO change detection task that contained four types of natural disasters: fires, floods, hurricanes and landslides. We benchmarked our model against a state-of-the-art AI model and achieved better results for floods and comparable results in the remaining three categories, without any training. We found that the PANN was able to achieve these results due to its ability to extract meaningful features from the inputs and sort them accordingly in the feature space. Being able to navigate the feature space, therefore, plays a vital role in the model’s performance. The training-free nature of the PANN makes it an ideal candidate for use on-board satellites, processing EO data at the extreme edge.

### v2

@@ Abstract
Earth observations from low Earth orbit satellites provide vital information for decision makers to better manage time–sensitive events such as natural disasters. For the data to be most effective for first responders, low latency is required between data capture and its arrival to decision makers. A major bottleneck is in the bandwidth–limited downlinking of the data from satellites to ground stations. One approach to overcome this challenge is to process at least some of the data on-board and prioritise pertinent data to be downlinked. In this study we propose a Physics Aware Neuromorphic Network (PANN) to detect changes caused by natural disasters from a sequence of multi-spectral satellite images and produce a change map, enabling relevant data to be prioritised for downlinking. The PANN used in this study is motivated by physical neural networks comprised of nano-electronic circuit elements known as “memristors” (nonlinear resistors with memory). The weights in the network are dynamic and update in response to varying input signals according to memristor equations of state and electrical circuit conservation laws. The PANN thus generates physics–constrained dynamical output features which are used to detect changes in a natural disaster detection task by applying a distance-based metric. Importantly, this makes the whole model training–free, allowing it to be implemented with minimal computing resources. The PANN was benchmarked against a state-of-the-art AI model and achieved comparable or better results in each natural disaster category. It thus presents a promising solution to the challenge of resource–constrained on-board processing.
@@ Introduction
Traditional algorithms have been developed to detect changes in EO images, for example BFAST [13], which has been applied to detecting changes in forests [14]. Algorithms such as BFAST rely on spectral indices as input. These are a relatively simple pixel comparison method to highlight objects of interest such as burn areas [15], diseased crops [16] and flood mapping [17]. These indices are effective as they utilise domain knowledge regarding the object of interest. For example, the normalised difference vegetation index is commonly used to incorporate knowledge about the wavelengths that the Chlorophyll pigment absorbs [18]. More sophisticated indices, which involve more than just a normalised difference between two spectral bands, have been proposed to improve the detection [15, 19]. The advantage of spectral indices is that they provide a simple, low computational cost method to determine objects of interest quickly, with reasonable accuracy. Additionally, they do not require training or large datasets to deploy, unlike ML models.
@@
AI models based on deep ML, specifically Convolutional Neural Networks (CNNs), have been applied to several natural disaster tasks using EO data, including flood segmentation [20, 21], detecting earthquake affected buildings [22] and volcanic eruptions [23]. To increase the performance of these AI models, often the number of layers and weights within the model needs to be increased. Chintalapati et al. [24] showed how the architecture of CNN models has evolved to increase performance on an EO image classification task, which resulted in an increase in the number of weights in the models, with the two best performing models, CoCa [25] and BASIC-L [26], using 2.1 and 2.44 billion parameters, respectively. It was noted that most deep learning AI models have so many parameters that they would exceed the memory budget on most EO satellites. In recent years, work has been done to design smaller AI models specifically for use on-board satellites [27, 28]. Furthermore, missions such as $\Phi$-Sat-1[29] and OPS-SAT [30] have also designed satellites to accommodate AI models on-board. Although these models have been specifically designed for on-board use, all the training still occurs on the ground because of the large computational overhead. Only inference is performed on-board. Since AI models are trained on the ground, they often cannot be trained with the on-board sensor data, leading to degraded performance [31, 32]. Work has been done to retrain models, once operational, with new data as it is acquired [33, 34].
@@
In this study, we introduce the PANN, a AI model with a novel architecture modelled after physical neural networks comprised of nano-electronic circuit elements known as memristors (nonlinear resistors with memory) that mimic synapses [35, 36, 37]. Like biological synapses, the memristive network weights continuously update in response to varying input signals, constrained by physical laws embedded within the model. The PANN, therefore, actively learns as new inputs are fed into the network, such as from a sensor on-board a satellite. As such, the PANN takes a significant shift away from the conventional artificial neural network AI paradigm, as the gradient–descent, back–propagation training of weights is completely replaced by physics–constrained dynamical equations. This is also in contrast to physics–informed neural networks, where the physics is embedded as a component of the model in the artificial neural network paradigm. Typically, this is done by modifying activation functions, gradient descent optimisation techniques, network structure, or loss functions [38, 39]. The added physics normally takes the form of a partial differential equation and is task specific, for example, incompressible Navier-Stokes equations for fluid flow problems [40], whereas with the PANN, the physics is independent of the task.
@@
In this study, we use a natural disaster dataset released by Růžička et al. [28] to evaluate the performance of the PANN model on a change detection problem, that contains a time series of multi-spectral EO images from Sentinel-2. We benchmark the PANN against a state-of-the-art AI model and achieve comparable or better results. We also evaluate the performance of the PANN on a secondary dataset of the same natural disasters using multi-spectral EO images from LandSat-8, to compare the models performace on different sensors. Additionally, we use a distance measurement to detect changes, bypassing the need for an output layer altogether. Hence, we do not need to train any part of the model and the PANN does not need to be trained on the ground prior to deployment.
@@ Results
The PANN model is used here to detect regions of change from natural disasters using EO multi-spectral images. The change detection task requires the model to take in a sequence of five images, where the final image is after a natural disaster event. The model outputs a change feature map that highlights the regions affected by the natural disaster, by breaking the images into smaller tiles and assigning a change likelihood score for each tile. The PANN model used in this study is an adaptive network, with weights that adjust with each new input. This allows the PANN to generate features from the inputs without requiring any training (see Methods for details). To determine if a tile has changed, the distance between inputs are compared in the feature space rather than the input space (pixel comparison). By applying a distance-based metric directly to the feature space, no training is performed at any stage in the model. The dataset includes four different categories of natural disaster events: fires, floods, hurricanes and landslides. Additionally, each natural disaster event is accompanied by a change mask which is used only to evaluate the performance of the PANN model.
@@
Figure 1 shows the output change maps produced by the PANN model for each type of natural disaster event. Alongside the change maps are the RGB images directly before and after the natural disaster events and the associated change mask, which also includes clouds from the images directly before and after the natural disaster event. Although only one image is shown before the natural disaster event in Fig. 1, a total of four images before the natural disaster event were used to create the displayed change maps. For the four natural disaster events, the main areas of change have been correctly identified as changes in the change maps, with a limited amount of noise in the images being considered as a change, such as the vegetation colour change in the fire natural disaster event (top row), or the sediment being washed into the ocean in the flood natural disaster event (second row). Clouds are detected as changes, but are masked out of the change maps shown in Fig. 1, since they are ignored when evaluating the change maps to allow for a more direct comparison to a state-of-the-art deep learning model (see Methods for details).
@@
Figure 1: RGB images (columns 1 and 2) directly before and after a natural disaster event, the corresponding target change mask (column 3) and the output change map produced by the PANN model (column 4) for each type of natural disaster: Top row – an example fire natural disaster event; Second row – an example flood natural disaster event; Third row – an example hurricane natural disaster event; Bottom row – an example landslide natural disaster event. Cloud-cover has been masked from the images and corresponds to yellow in the change mask, with genuine differences coloured aqua. Satellite images are from Sentinel-2 complied by Růžička et al. and available in their git[57].
@@
For each natural disaster event a change map was created, with the order in which the natural disasters were passed into the PANN model being randomised. We evaluated the quality of the change maps by calculating the Area Under the Precision—Recall Curve (AUPRC), which we create tile–wise. The change likelihood score for each tile is given by equation (3) in Methods, which depends on the distance metric used to measure the position of the tiles in the feature space. Table 1 reports the overall AUPRC as a percentage for each of the natural disaster categories using three different distance metrics, Euclidean, Cosine and Correlation distances. The Correlation distance significantly outperforms the other distance metrics and is also much more consistent across the different natural disaster classes.
@@
[…]
@@
[…]
To assess how well the PANN model generalises to data from different sensors, we evaluated the model on a subset of the natural disaster events using EO images from LandSat-8 and compared the results to the Sentinel-2 data (see Methods for details). Table 3 reports the performance of the PANN and baseline models for the LandSat-8 data and the comparative results for the equivalent subset in the Sentinel-2 data. From the results it can be seen that the PANN model outperforms the baseline method for both the data from the Sentinel-2 satellites and the LandSat-8 satellite for every natural disaster category. Both models achieve higher performance when using data from Sentinel-2 compared to LandSat-8. Additionally, the difference in performance between datasets is much larger for the baseline methods, which shows a significant decrease in performance when using the LandSat-8 data.
Model Fires Floods Landslides
S2 L8 S2 L8 S2 L8
Baseline 79.0 59.1 64.6 38.8 68.9 34.4
PANN $90.26\pm 0.02$ $86.07\pm 0.02$ $66.38\pm 0.03$ $53.9\pm 0.1$ $75.88\pm 0.02$ $53.3\pm 0.2$
[…]
@@
The AI models RaVAEn and PANN outperform the baseline method, since they can extract the most relevant features from the tiles and group similar tiles together, while keeping dissimilar tiles further apart in their feature spaces. To visualise the high dimensional feature space of the PANN, we used the Uniform Manifold Approximation and Projection (UMAP) [59] method to reduce the dimensions of the feature space into 2D. The UMAP used 20 neighbours, a minimum distance of 0.4 and the correlation distance, unless otherwise stated. Figure 2 shows the UMAP 2D projection of the feature space for a flooding event. The left-hand side shows the position of the tiles in the image directly before and after the flooding event. For tiles after the flooding event, the percentage change in each tile is denoted by the colour bar. The right-hand side of the figure shows the same projection of the feature space but with the tile images instead of points. The change tiles are grouped together at the bottom of the projections and the remaining tiles are also sorted according to the features in the tiles, with greener farmland tiles on the left-hand side and tiles with the white objects on the right. Tiles that have missing values, from the north alignment of the swath, are also grouped towards the bottom right, with the tiles that have the missing values across the top being separated from the tiles with the missing values on the right-side of the tile. For comparison with the UMAP projection for the pixel space, see Supplementary Fig. S1 online.
@@ Discussion
In this study, we introduced the PANN model for a natural disaster change detection task and compared it against the state-of-the-art AI model RaVAEn and a baseline method. All three models receive the same multi-spectral EO images for the inputs and use the same change detection equation given by equation (3), for each tile in the image. The difference in the model performances, therefore, is attributed to their ability to extract the most relevant features from the inputs, while ignoring noise within the inputs. We note, however, that while the PANN model receives the same inputs as the other models, only the inputs corresponding to the bands selected by the automatic feature engineering process, on a scene by scene basis, are passed to the relevant individual networks. This inherently allows the PANN model to ignore irrelevant information as determined by the domain knowledge embedded in the feature engineering process. The baseline method does not extract any features from the inputs and instead does the comparison directly in the pixel space. This makes the baseline method more sensitive to noise in the images and, as expected, it exhibits the lowest performance as reported in Table 2. The AI models, on the other hand, extract meaningful features from the tiles and hence outperform the baseline method in every natural disaster category. The PANN and RaVAEn models achieve a comparable result in three of four of the natural disaster categories, demonstrating that both models extract equally useful features for these categories. For flood natural disasters, however, the PANN model outperforms the RaVAEn model and thus arguably extracts more meaningful features.
@@
Although being able to extract the features is the most important task for the PANN model, how the feature space is navigated is also important for the performance on the PANN model. This is seen in Table 1 where the same feature space can produce significantly different results based on the distance metric used to navigate the feature space. This is visualised in the UMAP projections in Fig. 3, where different distance metrics are used to create the projections. The feature space produced by the PANN has a high dimensionality, as such the relative Euclidean distance between points tends to decrease[60]. This would result in the change likelihood score decreasing, particularly for change tiles and is likely why the Euclidean distance does not produce the best results. What is surprising is the large difference between the Cosine and Correlation distances, given that both metrics are angle-based measures. The Cosine distance is given by equation (4) and measures the angle between the vectors in the feature space. The Correlation distance is the same as the Cosine distance except each sample is mean centred first, as given by equation (5). The mean-correction for the Correlation distance effectively changes the location of the origin and hence the angle between the vectors also changes [61]. One possible reason why the mean correction has such a large impact on the performance is that it helps reduce covariate shift in the output sequence that can arise from the dynamical nature of the PANN. The covariate shift effect can be seen in Fig. 3 (and Supplementary Fig. S2), where there is a much greater overlap of points at the different timesteps (that are not changes) in the Correlation projection compared to that of the Cosine projection. Interestingly, batch normalisation layers were introduced to deep neural networks to help reduce the covariate shift between layers in the networks [62] and have been widely accepted to speed up training times and increase performance. We note that the PANN does not have any batch normalisation layers, while the RaVAEn model does, and that the best performing metrics for the models are the Correlation and Cosine distance, respectively.
Figure 4: Comparison between the change maps produce by the PANN model using the Sentinel-2 (S2) and LandSat-8 (L8) dataset, for the same landslide natural disaster event. The corresponding change masks are also included.
From the results in Table 3, two clear trends are evident. The first is that both the PANN and baseline methods perform better when using data from Sentinel-2 than LandSat-8 across all natural disaster categories. This can largely be explained by the decrease in resolution. The Sentinel-2 data has a spatial resolution of 10 m, with a tile size $32\times 32$ pixels and a tile resolution of 320 m. While the LandSat-8 data has a spatial resolution of 30 m, with a tile size of $16\times 16$ pixels and tile resolution 480 m. As a result, the models are not able to resolve finer details when determining if a change has occurred in a LandSat-8 tile. This difference in resolution can be seen in Fig. 4, which shows the change mask and output change map for both datasets for one of the landslides. Additionally, in Fig. 4, more noise is evident in the LandSat-8 change map, compared to the Sentinel-2 change map, particularly in the areas where there is no change. This would also contributes to the decrease in performance. The second trend that can be clearly seen in Table 3, is that the decrease in performance using the LandSat-8 data relative to the Sentinel-2 data is significantly less for the PANN model compared to the baseline method. This shows that the PANN model is able to generalise to different sensor data much better than the baseline method, further showing that the PANN model is able to extract meaningful features from the input data.
@@
Our results show that the PANN model achieves comparable-to-better results than the RaVAEn state-of-the-art AI model for the natural disaster change detection task conducted in this study. Most importantly, the PANN was able to achieve this result by creating a mapping from the input space to the feature space without any training. The PANN is training-free as it continuously adapts with each new input. The model equations are relatively light–weight and thus all the processing of the data by the PANN model was able to be carried out using only CPUs in this study. The compute resources, time and memory, were measured for the PANN to process a patch sequence of $574\times 509$ pixels over 5 frames, corresponding to roughly a 25 km2 region. Using 8 cores on an Intel Xeon Scalable processor, the PANN model takes 19 s and uses approximately 1 GB of memory (see Supplementary Fig. S3). We note that the model could be optimised to further decrease memory requirements by reducing the floating-point precision of the PANN from 64 bit to 32 or 16 bit precision. We also note that the processing time for the PANN model is not directly comparable to the inference time of RaVAEn or other AI change detection models, as the processing time to evaluate a scene also includes updating of the PANN weights with each input. Unlike AI models, the PANN model does not require any additional time and compute resources for training. The physics–based equations that constitute the PANN model are a significant deviation from conventional deep multi-layered bipartite neural networks that typically use mathematical activation functions to perform nonlinear transformations and stochastic gradient decent with backpropagation as the learning algorithm to train networks weights into an optimal static state. As such, the iterative training process is computationally expensive and usually requires cloud-based GPUs [63], to produce a trained model. As the PANN simulates a physical device, based on memristive nanowire networks, we expect the computational resources to significantly decrease when tasks are implemented on the physical hardware (see e.g. [41, 64, 44, 48, 51]).
The PANN model is relatively sensor agnostic since it does not need to be pre-trained with data from a specific sensor and the model design of having a separate network for each spectral band (see Methods for details), allows it to receive images with varying number of spectral bands. Having a separate PANN for each band also allows a simple training–free method to add domain knowledge to the model in the form of an automatic feature engineering mechanism that determines which spectral bands and hence PANNs are used (see Methods for details). In this study, only relatively minor adjustments to the feature engineering mechanism were required when changing the sensor data from the Sentinel-2 satellite to the LandSat-8 satellite. In addition, having a separate PANN for each band combined with the feature engineering means only a subset of the PANNs are required, further reducing the computational resources. Having multiple smaller networks, however, means that interactions between bands may not be fully captured, as compared to a single larger network, where information from all the bands is integrated while being processed. Alternatively, another model design that could be considered is multiple networks with a hierarchical structure[65].
@@
In this study, we have demonstrated the feasibility of the PANN for detecting changes from natural disasters within the context of their potential deployment on-board satellites. Based on these simulation results, future work would include testing a prototype of the hardware device, based on memristive nanowire networks, which the PANN is modelled after. Based on previous studies with such neuromorphic devices [41, 64, 44, 48, 51], it is expected that implementation on the low-power hardware device will substantially reduce processing times and energy consumption, and could be achieved using only the on-board compute. The methodology outlined in this study could be used when implementing in hardware with one key change: the physical network devices may need larger electrode grids than the $16\times 16$ grid used here, as additional electrodes would be needed to read out voltages from the larger physical networks. To date, devices with up to 128 electrodes have been fabricated [66], although interestingly, experimental results suggest that using readouts from only a subset of electrodes may be sufficient to capture the most salient output features [41]. This may be attributed to how the electrodes integrate signals from many nanowires they make electrical contact with. This means that when implementing in a physical device, the option of using one large network that processes all the spectral bands together may be viable. Multiple configurations for the electrode placement would be possible, provided the contact electrodes evenly sample the underlying nanowire network.
In conclusion, we demonstrated a training-free AI model, the PANN, for a natural disaster EO change detection task that contained four types of natural disasters: fires, floods, hurricanes and landslides. We benchmarked our model against a state-of-the-art AI model and achieved better results for floods and comparable results in the remaining three categories, without any training. We also showed that the PANN model can generalise to other EO sensors. We found that the PANN was able to achieve these results due to its ability to extract meaningful features from the inputs and sort them accordingly in the feature space. Being able to navigate the feature space, therefore, plays a vital role in the model’s performance. The training-free nature of the PANN makes it an ideal candidate for use on-board satellites, processing EO data at the extreme edge.